OpenStudy (anonymous):

need help solving for a: 5a+3b=10

OpenStudy (angela210793):

5a=10-3b a=(10-3b)/5

OpenStudy (anonymous):

so I flop the ten and divide through with 5?

OpenStudy (angela210793):

yes :)

OpenStudy (anonymous):

ty

OpenStudy (angela210793):

yw :)

OpenStudy (anonymous):

so a=2-3/5b?